2. Find all eigenvalues and eigenvectors of the matrix 3 2 3 4
Answer #1

Let A= 3 2 3 4 | A-Il=0 2 3-) 3 4-7 → (3-4) (4-1)-6=0 7 12-70+8²-6=0 72_70+6=0 → X-6X-+6=0 > > (4-6) -(2-6)=0 = x= 1,6 1 andLet x = be X2 an eigenvector of a corresponding to x=6. Then (A-61)x=0 -3 74 = - 3x + 2x2=0 , 3x-2x=0 - 3x = 2x2 Let x=2. The